OpenStudy (anonymous):

Simplify: (3x+2y)(3x-2y)= 9x2-4y2 9x2+4y2 9x2+12xy+4y2 None of the above

OpenStudy (anonymous):

you would have to know that \[(a+b)(a-b)=a^2-b^2\]

OpenStudy (anonymous):

I got C

OpenStudy (anonymous):

no...that is not it.... look at what i have given you ^^, realize you have almost the same thing with your a=3x and your b=2y

OpenStudy (anonymous):

do you see that??

OpenStudy (anonymous):

so using what i have written above, \[(3x+2y)(3x-2y)=(3x)^2-(2y)^2\]

OpenStudy (anonymous):

what do you get when you simplify that??

OpenStudy (anonymous):

I SEE THAT I SEE THT OHHHHHHHH I SEEEEE THAT

OpenStudy (anonymous):

9x2+4y2